A 12-by-12 inches ingot iron casting 60-in high is stripped from its mold while its temperature is 1740 °F. If it stands on end on the floor of a large foundry having wall, floor, and roof temperatures of 80°F, what is the rate of radiant-heat exchange between the casting and the roof wall, and floor surfaces?
1
Expert's answer
2021-10-27T07:02:14-0400

Rate of radiant heat is given by

"\frac{Q}{t}=\sigma\epsilon AT^4"

Where "\sigma" =5.67*10-8J/sm2K4

A=12*12=144

"\epsilon=1"

"\frac{Q}{t}=\sigma\epsilon AT^4"

"=5.67*10^{-8}*1*144*80^4"

=334.43kJ/s